You added 20g of compound Cin 125 mL water. The MW of the compound C is 285g/mol. What the final molarity of the compound in that 125mL water?


1
Expert's answer
2021-08-24T02:19:38-0400

СM = n/V

n = m/M

n = 20/285 = 0.07 g/mol

CM = 0.07/0.125 = 0.56 mol/L